# Bias in Language Models: Beyond Trick Tests and Towards RUTEd Evaluation

### v2

@@ Abstract
Standard benchmarks of bias and fairness in large language models (LLMs) measure the association between social attributes implied in user prompts and short LLM responses. In the commonly studied domain of gender-occupation bias, we test whether these benchmarks are robust to lengthening the LLM responses as a measure of Realistic Use and Tangible Effects (i.e., RUTEd evaluations). From the current literature, we adapt three standard bias metrics (neutrality, skew, and stereotype), and we develop analogous RUTEd evaluations from three contexts of real-world use: children’s bedtime stories, user personas, and English language learning exercises. We find that standard bias metrics have no significant correlation with the more realistic bias metrics. For example, selecting the least biased model based on the standard “trick tests” coincides with selecting the least biased model as measured in more realistic use no more than random chance. We suggest that there is not yet evidence to justify standard benchmarks as reliable proxies of real-world biases, and we encourage further development of context-specific RUTEd evaluations.
@@
Kristian Lum1, Jacy Reese Anthis1,2, Kevin Robinson3, Chirag Nagpal3, Alexander D’Amour4 1University of Chicago, 2Stanford University, 3Google Research, 4Google DeepMind
@@ 1 Introduction
We argue that standard benchmarks constitute “trick tests”: decontextualized evaluations based on contrived scenarios designed to elicit a simplified correlation between model output and a sensitive attribute rather than as best estimates of the real-world effects of model use. We contrast these tests with novel evaluations that are grounded, at least to some extent, in Realistic Use and Tangible Effects, or RUTEd (“rooted”) evaluations. The need for RUTEd evaluations echoes calls for sociotechnical evaluations of ML systems, beyond the current focus on “a small space of the potential harm” (Weidinger et al., 2023). We conduct this study in the context of gender-occupation bias, the most common association tested in bias benchmarks (Weidinger et al., 2023), and this association allows us to sidestep much of the subjectivity and debates around other social contexts, such as race and socioeconomic status (Blodgett et al., 2021).
@@ 2.1 Static word embeddings
Popular intrinsic bias metrics of this sort include the Word Embedding Association Test (WEAT) benchmark Caliskan et al. (2017)) and the similar approach of Bolukbasi et al. (2016). Both aggregate cosine similarity measures between words associated with different identity groups (e.g., “he,” “she”) with words in a domain of interest (e.g., occupations). In the paper that introduced the intrinsic-extrinsic dichotomy (Goldfarb-Tarrant et al., 2020), the intrinsic metric of WEAT (in both English and Spanish) was contrasted with extrinsic metrics based on models that used those embeddings for the tasks of coreference resolution and hate speech detection.
@@ 3 Evaluations
We adapt a standard benchmark from the literature,the Gender Sensitivity task from the popular BIG-bench testing suite (Srivastava et al., 2022), and compare it with three long-form text generation tasks based on realistic use cases. Each evaluation has three derived metrics, and we test nine LLMs. We focus exclusively on the context of gender-occupation bias associated with male and female genders. While there are many other types of bias that could arise, and we acknowledge the limitations of the binary gender framework, we chose this paradigm for internal consistency across evaluations and consistency with prior work. Gender-occupation is the most prevalent paradigm in the literature on bias in NLP with approximately 17% of bias evaluations using it according to a recent literature review (Weidinger et al., 2023).
@@ 3.2 RUTEd evaluation tasks
To test whether standard benchmarks measure a stable feature of LLMs that can be directly translated to real-world contexts, we drew on media discussions to identify common use cases that are suitable for evaluation and more connected to tangible effects. Note that our evaluation tasks are only intended for research, not deployment as real-world evaluations. Example outputs for each task are provided in Appendix C.
@@
Public data on LLM use is limited, but one example that is frequently discussed in the media is generating bedtime stories for small children (BedtimeStory.ai, 2023; Kobie, 2023; McGuinness, 2023; OpenAI, 2023; Srivastava, 2023). This is reportedly a frequent use case in which models perform relatively well, and bedtime stories are a daily interaction for many parents. Storytelling has the ability to spark a child’s imagination and shape what they think of as possible. For this reason, we believe that reinforcing stereotypes—particularly in the most common bias example of gender-stereotyped occupations—may directly lead to real-world harm. It might subtly influence children’s beliefs about the types of occupations available to them.
@@
We prompt the model to generate sentences that could be used in instruction to demonstrate the use of pronouns to a language learner in the category of English as a Second Language (ESL). In this case, a significant gender imbalance could have similar harms as those of bedtime stories for children or for adults who are getting to know the culture and social norms of English-speaking regions, as well as harm through language learning as students learn to express different concepts at different rates and learn to favor certain words in certain contexts. To generate the sentences, we input the prompt:
@@ 3.3 Metrics
For the long-form text generation tasks, we must statistically estimate probabilities. For the Llama-2, GPT-4, and Mixtral-8x7B models, we generate $n=30$ replicates per task and occupation; for the Flan-PaLM models, we generate $n=64$. Models were set to default temperature with no minimum token probability and with the aforementioned maximum tokens for each context. Then, for each occupation, $o$, we calculate the proportion of replicates for that gender in which the generated text was about males, $\hat{p}_{o}^{m}$, and females, $\hat{p}_{p}^{f}$. Those for which greater than half of the pronouns refer to males are categorized as “male” replicates; the others are categorized as “female.” Because each occupation has an associated gender-majority, we also calculate the proportion of replicates that were gender-stereotypical, $\hat{p}_{o}^{s}$, and gender anti-stereotypical, $\hat{p}_{o}^{a}$. Replicates with no such pronouns are dropped. For clarity, we define metrics with the hatless notation and plug in $\hat{p}$ when necessary.
@@
Rather than the absolute difference from parity, we define the skew metric as the average tendency of the model to return male words or replicates instead of female words or replicates. $m^{\text{skew}}=\frac{1}{O}\sum_{o}\left(p^{\text{m}}_{o}-p^{\text{f}}_{o}\right)$. If male words/replicates have a higher probability systematically across all considered occupations, this metric is positive. Conversely, if female words/replicates have a systematically higher probability, the metric is negative.
@@
Some studies have measured the difference between stereotypical and anti-stereotypical token generation (e.g., de Vassimon Manela et al., 2021; Nadeem et al., 2020). We define the stereotype metric analogously to neutrality: $m^{\text{stereotype}}=\frac{1}{O}\sum_{o}\left(p_{o}^{s}-p_{o}^{a}\right)$. To create a standard benchmark evaluation, the sum ranges over occupations in the benchmark. Here, $O$ represents the number of occupations used in the task.
@@ 3.4 Statistical uncertainty
Because the probabilities in the decontextualized evaluations are directly observed, there is no statistical uncertainty (and therefore no error bars in Figure 1). For the RUTEd evaluations, we estimate the sampling variance of each estimated probability. To calculate this, we first note that one component used in each metrics is $d_{o}=\hat{p}_{o}^{m}-\hat{p}_{o}^{f}=\hat{p}_{o}^{m}-(1-\hat{p}_{o}^{m})$. For the RUTEd tasks, using a simple plug-in estimator of the sampling variance of $\hat{p}_{o}^{m}$, we get $d_{o}=2\hat{p}_{o}^{m}-1$ with sampling variance $\hat{\sigma}_{o}^{2}=4\frac{\hat{p}_{o}^{m}(1-\hat{p}_{o}^{m})}{n}$. We apply this to each of the three metrics in the following two sections—where the variance of skew and stereotype are equal, and therefore only one derivation is needed.
@@ 3.5 Models
We generated content and calculated metrics for models from four different families: Llama-2, Flan-PaLM, GPT-4-0125-preview, and Mixtral-8x7B. For the Llama-2 and Flan-PaLM models that have base models (i.e., only pre-trained) and chat (i.e., pre-trained and instruction-tuned) versions, we used the chat versions to mimic consumer use. For Llama-2 and Flan-PaLM, we evaluated several model sizes: the Llama-2 7, 13, and 70 billion parameter models (Touvron et al., 2023), and for Flan-PaLM, we evaluate the extra-small (XS), small (S), medium (M), and large (L) models (Chung et al., 2022). For GPT-4 and Mixtral-8x7B, we did not have access to next-word probabilities from the model providers, so we could not run the decontextualized standard benchmarks for these models, and therefore these models only contribute to the between-RUTEd evaluations analysis.
@@ 4.1 Correlations between standard benchmarks and RUTEd evaluations
A decision-maker may instead be selecting across all the models. For neutrality, the least biased is still Llama-2 13B. For skew, the least biased is Flan-PaLM L; note that for skew and stereotype, the values can be negative, and if some are, then still the lowest score (i.e., most negative score) is considered the least biased. For stereotype, the least biased is Flan-PaLM XL. Among the nine RUTEd evaluations, none of them assert the same as the corresponding standard benchmark. If we selected models at random, we would be correct approximately one in seven times, as we are excluding GPT-4 and Mixtral for this comparison.
@@ 4.2 Correlations between RUTEd evaluations
The inconsistency is clearer in the three pairwise correlations averaged across metrics, shown in Table 2, which shows Spearman’s rank correlation between each pair of RUTEd evaluations, averaged over the three potential metrics of interest. While each correlation is positive, they are near zero. This suggests that, in these three contexts, selecting or ranking models based on one context would not be a reliable way to identify the least biased models for application to another context. This echoes arguments for context-specific fairness from perspectives such as statistical theory (Anthis and Veitch, 2023), inverse reinforcement learning (Blandin and Kash, 2024), and social computing (Madaio et al., 2022).
@@ 4.4 Mode collapse
Mode collapse, a phenomenon in which a generative model produces only very similar outputs (Salimans et al., 2016), could distort bias estimates if the same replicate—possibly with small variation in wording—is generated repeatedly. We analyzed the 10,800 replicates for Llama-2 models (3 models, 3 RUTEd evaluations, 40 occupations, and 30 replicates per group) by first finding the groups of 30 replicates with the same model, evaluation, and occupation that had the highest average cosine similarity amongst themselves, using all-MiniLM-L6-v2 for sentence embeddings. We then manually inspected the groups with the most similarity and a random sample of other groups. We find a variety of replicates, even within the groups with the most cross-replicate similarity, suggesting that our findings are not the result of mode collapse. Further work in this area could introduce additional prompt variations as well as variation in model temperature or other hyperparameters.
@@ 6 Limitations
Finally, we note that while our results suggest caution when using standard bias benchmarks in real-world application, they do not diminish the contributions or usefulness of these benchmarks or other prior work. The field of algorithmic fairness has built technical and empirical frameworks step-by-step, and shifting paradigms (e.g., towards decoder-only transformer architectures) have made it challenging to keep pace with real-world use.
@@ Appendix B Prompt variation
Table B3: The 10 “ESL Leaning Exercises” prompt templates used to test if results depend on particular prompt wording. Template 0 is the template used in the main analysis.

### v3

@@ Abstract
Standard benchmarks of bias and fairness in large language models (LLMs) measure the association between the user attributes stated or implied by a prompt and the LLM’s short text response, but human-AI interaction increasingly requires long-form and context-specific system output to solve real-world tasks. In the commonly studied domain of gender-occupation bias, we test whether these benchmarks are robust to lengthening the LLM responses as a measure of Realistic Use and Tangible Effects (i.e., RUTEd evaluations). From the current literature, we adapt three standard bias metrics (neutrality, skew, and stereotype) and develop analogous RUTEd evaluations from three contexts of real-world use: children’s bedtime stories, user personas, and English language learning exercises. We find that standard bias metrics have no significant correlation with the more realistic bias metrics. For example, selecting the least biased model based on the standard “trick tests” coincides with selecting the least biased model as measured in more realistic use no more than random chance. We suggest that there is not yet evidence to justify standard benchmarks as reliable proxies of real-world AI biases, and we encourage further development of evaluations grounded in particular contexts.
@@
Kristian Lum1,4, Jacy Reese Anthis1,2, Kevin Robinson3, Chirag Nagpal3, Alexander D’Amour4 1University of Chicago, 2Stanford University, 3Google Research, 4Google DeepMind
@@ 1 Introduction
We argue that standard benchmarks constitute “trick tests”: decontextualized evaluations based on contrived scenarios designed to elicit a simplified correlation between model output and a sensitive attribute rather than as best estimates of the real-world effects of model use. We contrast these tests with novel evaluations that are grounded, at least to some extent, in Realistic Use and Tangible Effects, or RUTEd (“rooted”) evaluations. The need for RUTEd evaluations echoes calls for sociotechnical evaluations of ML systems, beyond the current focus on “a small space of the potential harm” (Weidinger et al., 2023). We conduct this study in the context of gender-occupation bias, the most common association tested in bias benchmarks (Weidinger et al., 2023). In addition to the societal importance of this association, it allows us to sidestep much of the subjectivity and debates around other social contexts, such as race and socioeconomic status (Blodgett et al., 2021).
@@
In summary, we make the following contributions:
  • We review the evolution of NLP bias evaluations through static word embeddings, LLMs, and fine-tuning techniques. In particular, we highlight changes in the distinction between “intrinsic” and “extrinsic” evaluations.
  • We provide a new conceptual framework, RUTEd evaluations, that can include a variety of bias and fairness evaluations more applicable to real-world, general-purpose LLM use.
  • We derive three metrics from the extant literature (stereotype, neutrality, and skew) and compare them to three analogous metrics—each tested on three long-form text generation use cases (Bedtime Stories, User Personas, and ESL Learning Exercises).
  • Across nine popular LLMs and with several robustness checks, we show that standard benchmarks do not predict the RUTEd evaluations and that RUTEd evaluations do not predict each other, showing the need to move beyond standard benchmarks and incorporate social context.
@@ 2.1 Static word embeddings
Popular intrinsic bias metrics of this sort include the Word Embedding Association Test (WEAT) benchmark Caliskan et al. (2017) and the similar approach of Bolukbasi et al. (2016). Both aggregate cosine similarity measures between words associated with different identity groups (e.g., “he,” “she”) with words in a domain of interest (e.g., occupations). In the paper that introduced the intrinsic-extrinsic dichotomy (Goldfarb-Tarrant et al., 2020), the intrinsic metric of WEAT (in both English and Spanish) was contrasted with extrinsic metrics based on models that used those embeddings for the tasks of coreference resolution and hate speech detection.
@@ 3 Evaluations
We adapt a standard benchmark from the literature, the Gender Sensitivity task from the popular BIG-bench testing suite (Srivastava et al., 2022), and compare it with three long-form text generation tasks based on realistic use cases. We adopt this task because it is amenable to testing and comparison. Other benchmarks, such as StereoSet (Nadeem et al., 2020) and BBQ Parrish et al. (2022), have more idiosyncrasies. For example, the first item described in the StereoSet paper presents the model with the prompt ‘‘Girls tend to be more ____ than boys’’ and asks the model to fill in the blank with “soft,” “determined,” or “fish.” Lengthening this format seems to require more changes—or more degrees of freedom—such as transforming the prompt into an open-ended question and counting words related to the multiple-choice options. Such changes could also amplify the idiosyncrasies in the short-form benchmark, including the specific choices of multiple-choice options (e.g., “soft” may have many common related words, regardless of relevance).
Each evaluation has three derived metrics, and we test nine LLMs. We focus exclusively on the context of gender-occupation bias associated with male and female genders. While there are many other types of bias that could arise, and we acknowledge the limitations of the binary gender framework, we chose this paradigm for internal consistency across evaluations and consistency with prior work. Gender-occupation is the most prevalent paradigm in the literature on bias in NLP, present in approximately 17% of bias evaluations according to a recent literature review (Weidinger et al., 2023).
The decision to focus this study on gender-occupation limits the generalizability of the results. There may be types of bias for which decontextualized metrics provide robust information across use cases. However, by demonstrating that decontextualized bias metrics can—in the instance of gender-occupation bias—tell a misleading story about how bias manifests in more realistic settings, we show that decontextualized metrics are inadequate to guarantee “unbiased” performance in general.
@@ 3.2 RUTEd evaluation tasks
To test whether standard benchmarks measure a stable feature of LLMs that can be directly translated to real-world contexts, we drew on media discussions to identify common use cases that are suitable for evaluation and more connected to tangible effects. The RUTEd evaluation tasks were developed as a contrast to the standard benchmarks for research purposes. While we believe they more closely mirror realistic use, they are not validated for realism or use by practitioners to make real-world decisions. For example, an English language teacher should not rely exclusively on these to decide which LLM to use for content creation assistance. A teacher concerned about the gender distribution in examples they are creating would need to test specifically for their intended use. Example outputs for each task are provided in Appendix C.
@@
Public data on LLM use is limited, but one example that is frequently discussed in the media is generating bedtime stories for small children (BedtimeStory.ai, 2023; Kobie, 2023; McGuinness, 2023; OpenAI, 2023; Srivastava, 2023). This is reportedly a frequent use case in which models perform relatively well, and bedtime stories are a daily interaction for many parents. Storytelling has the ability to spark a child’s imagination and shape what they think of as possible. For this reason, we believe that reinforcing stereotypes—particularly in the most common bias example of gender-stereotyped occupations— may be an area of real world concern for model users, as it has the potential to subtly influence children’s beliefs about the types of occupations available to them.
@@
We prompt the model to generate sentences that could be used in instruction to demonstrate the use of pronouns to a language learner in the category of English as a Second Language (ESL). In this case, a significant gender imbalance could impact how adults who are getting to know the culture and social norms of English-speaking regions learn to express different concepts and learn to favor certain words in certain contexts. To generate the sentences, we input the prompt:
@@ 3.3 Metrics
For the long-form text generation tasks, we must statistically estimate probabilities. For the Llama-2, GPT-4, and Mixtral-8x7B models, we generate $n=30$ replicates per task and occupation; for the Flan-PaLM models, we generate $n=64$. Models were set to default temperature with no minimum token probability and with the aforementioned maximum tokens for each context. Then, for each occupation, $o$, we calculate the proportion of replicates for that gender in which the generated text was about males, $\hat{p}_{o}^{m}$, and females, $\hat{p}_{o}^{f}$. Those for which greater than half of the pronouns refer to males are categorized as “male” replicates; the others are categorized as “female.” Because each occupation has an associated gender-majority, we also calculate the proportion of replicates that were gender-stereotypical, $\hat{p}_{o}^{s}$, and gender anti-stereotypical, $\hat{p}_{o}^{a}$. Replicates with no such pronouns are dropped. For clarity, we define metrics with the hatless notation and plug in $\hat{p}$ when necessary.
@@
Rather than the absolute difference from parity, we define the skew metric as the average tendency of the model to return male words or replicates instead of female words or replicates. $m^{\text{skew}}=\frac{1}{O}\sum_{o}\left(p^{\text{m}}_{o}-p^{\text{f}}_{o}\right)$. If male words/replicates have a higher probability systematically across all considered occupations, this metric is positive. Conversely, if female words/replicates have a systematically higher probability, the metric is negative. This metric addresses the tendency of the model towards male or female outputs, irrespective of the current gender distribution in a profession.
@@
Some studies have measured the difference between stereotypical and anti-stereotypical token generation (e.g., de Vassimon Manela et al., 2021; Nadeem et al., 2020). We define this metric: $m^{\text{stereotype}}=\frac{1}{O}\sum_{o}\left(p_{o}^{s}-p_{o}^{a}\right)$. To create a standard benchmark evaluation, the sum ranges over occupations in the benchmark. Here, $O$ represents the number of occupations used in the task. Positive values of this metric indicate that generations were more likely to conform with stereotypes.
@@ 3.4 Statistical uncertainty
Because the probabilities in the decontextualized evaluations are directly observed, there is no statistical uncertainty (and therefore no error bars in Figure 1). For the RUTEd evaluations, we estimate the sampling variance of each estimated probability. To calculate this, we first note that one component used in each metric is $d_{o}=\hat{p}_{o}^{m}-\hat{p}_{o}^{f}=\hat{p}_{o}^{m}-(1-\hat{p}_{o}^{m})$. For the RUTEd tasks, using a simple plug-in estimator of the sampling variance of $\hat{p}_{o}^{m}$, we get $d_{o}=2\hat{p}_{o}^{m}-1$ with sampling variance $\hat{\sigma}_{o}^{2}=4\frac{\hat{p}_{o}^{m}(1-\hat{p}_{o}^{m})}{n}$. We apply this to each of the three metrics in the following two sections—where variance of skew and stereotype are equal, and only one derivation is needed.
@@ 3.5 Models
We generated content and calculated metrics for models from four different families: Llama-2,  [^1] Llama-2 evaluations were run on the University of Chicago cluster, prior to author KL’s affiliation with Google DeepMind. Flan-PaLM, GPT-4-0125-preview, and Mixtral-8x7B. For the Llama-2 and Flan-PaLM models that have base models (i.e., only pre-trained) and chat (i.e., pre-trained and instruction-tuned) versions, we used the chat versions to mimic consumer use. For Llama-2 and Flan-PaLM, we evaluated several model sizes: the Llama-2 7, 13, and 70 billion parameter models (Touvron et al., 2023), and for Flan-PaLM, we evaluate the extra-small (XS), small (S), medium (M), and large (L) models (Chung et al., 2022). For GPT-4 and Mixtral-8x7B, we did not have access to next-word probabilities from the model providers, so we could not run the decontextualized standard benchmarks for these models, and therefore these models only contribute to the between-RUTEd evaluations analysis.
@@ 4.1 Correlations between standard benchmarks and RUTEd evaluations
A decision-maker may instead be selecting across all the models. For neutrality, the least biased is still Llama-2 13B. For skew, the least biased is Flan-PaLM L; note that for skew and stereotype, the values can be negative, and if some are, then still the lowest score (i.e., most negative score) is considered the least biased. For stereotype, the least biased is Flan-PaLM M. Among the nine RUTEd evaluations, none of them assert the same as the corresponding standard benchmark. If we selected models at random, we would be correct approximately one in seven times, as we are excluding GPT-4 and Mixtral for this comparison.
@@ 4.2 Correlations between RUTEd evaluations
The inconsistency is clearer in the three pairwise correlations averaged across metrics, shown in Table 2, which shows Spearman’s rank correlation between each pair of RUTEd evaluations, averaged over the three potential metrics of interest. While each correlation is positive, they are near zero. This suggests that selecting or ranking models based on one context would not be a reliable way to identify the least biased models for application to another context. This echoes arguments for context-specific fairness from perspectives such as statistical theory (Anthis and Veitch, 2023), inverse reinforcement learning (Blandin and Kash, 2024), and social computing (Madaio et al., 2022).
@@ 4.4 Mode collapse
Mode collapse, a phenomenon in which a generative model produces only very similar outputs (Salimans et al., 2016), could distort bias estimates if the same replicate—possibly with small variation in wording—is generated repeatedly. We analyzed the 10,800 replicates for Llama-2 models (3 models, 3 RUTEd evaluations, 40 occupations, and 30 replicates per group) by first finding the groups of 30 replicates with the same model, evaluation, and occupation that had the highest average cosine similarity amongst themselves, using all-MiniLM-L6-v2 for sentence embeddings. We manually inspected the groups with the most similarity and a random sample of other groups. We find a variety of replicates, even within the groups with the most cross-replicate similarity, suggesting that our findings are not the result of mode collapse. Future work could vary temperature or other hyperparameters.
@@ 6 Limitations
Finally, we note that while our results suggest caution when using standard bias benchmarks in real-world application, they do not diminish the contributions or usefulness of these benchmarks or other prior work. The field of algorithmic fairness has built technical and empirical frameworks step by step, and this has been especially challenging as model architectures have evolved, such as the shift towards decoder-only transformer architectures.
@@ Appendix B Prompt variation
Table B3: The 10 “ESL Learning Exercises” prompt templates used to test if results depend on particular prompt wording. Template 0 is the template used in the main analysis.